How much of a 45% solution must be added to 1 liter of a 60% solution in order to make a 50% solution?
Answer by tommyt3rd(5050)

0.45x+0.60*1=(x+1)*0.5
0.6-0.5=0.5x-0.45x
1.0=0.05x
x=20 liters